# Solving the Poisson equation using coupled Markov chains

### v3

@@ Appendix E Verifying the assumptions in the AR(1) case
Algorithm E.1 Simulating pairs of ${\rm AR}(1,\phi,\sigma^{2})$ chains with reflection-maximal coupling.
@@
Lemma 61.
For Algorithm E.1, the meeting time $\tau=\inf\{n\geq 1:X_{n}=Y_{n}\}$ satisfies
@@
We consider the following equivalent construction of $X$ and $Y$ as in Algorithm E.1. Let $(W_{n})$ be a sequence of independent $\text{Normal}(0,1)$ random variables. Set $X_{0}=x_{0}$ and define
@@
Combining Lemma 61 with Proposition 62, we obtain the following.
Corollary 63.
For Algorithm E.1, the meeting time $\tau=\inf\{n\geq 1:X_{n}=Y_{n}\}$ satisfies
@@
where with the same constants as in Proposition 62, $\tilde{C}(x,y)=\frac{2}{\tilde{\beta}}+|\frac{x-y}{2\sigma}|+3$ and $\bar{\beta}=\tilde{\beta}^{\frac{\log(\phi)}{\log(\tilde{\beta})+\log(\phi)}}$, and satisfies $\pi\otimes\pi(\tilde{C})<\infty$.
@@
We conclude by Lemma 64. ∎
In combination with Lemma 61, we obtain the following.
Corollary 66.
For Algorithm E.1, the meeting time $\tau=\inf\{n\geq 1:X_{n}=Y_{n}\}$ satisfies
@@
where $\tilde{\beta}\in(0,1)$ is as in Proposition 65, and $\tilde{C}(x,y)=1+(1-\phi^{2})\left(\frac{x_{0}-y_{0}}{2\sigma}\right)^{2}$ satisfies $\pi\otimes\pi(\tilde{C})<\infty$.
@@
where $\gamma=\beta^{\frac{\log(\alpha)}{\log(\beta)+\log(\alpha)}}$ and $\mathbb{P}_{\nu}(\tau>n)$ is as in Lemma 67.
@@
Lemma 67 provides that
@@
where we have used Devroye et al. (2018, Theorem 1.3) in the first line. From Proposition 65, we find that
@@
we may deduce that $\mathbb{P}_{\mu}(\tau>n)\leq 2\tilde{\beta}^{n}$. Hence, we obtain by Corollary 68,
@@
Appendix F Hamiltonian Monte Carlo
F.1 Basic version of HMC
Consider a target distribution on $\mathbb{R}^{d}$ with density $\pi$ and gradient of the log-density $\nabla\log\pi$, that can be evaluated pointwise. With Hamiltonian Monte Carlo (e.g. Neal 2011), we augment the state space $\mathbb{R}^{d}$ with a momentum variable, also taking values in $\mathbb{R}^{d}$. The momentum is periodically resampled from a $\text{Normal}(0,M)$ distribution, with covariance matrix $M\in\mathbb{R}^{d\times d}$. Algorithm F.1 describes one step of the leapfrog integrator, which updates a pair $(x,m)$ of position and momentum variables in a way that mimics certain Hamiltonian dynamics.
@@ F.1 Basic version of HMC
Algorithm F.1 Function $(x,m)\mapsto\text{leapfrog}(x,m)$, with stepsize $\epsilon>0$.
The transition kernel of HMC is described in Algorithm F.2. From an initial position, it draws an independent momentum from a centered multivariate Normal, then performs $L$ leapfrog steps, and finally the terminal position is accepted or not as the new state, with a Metropolis–Rosenbluth–Teller–Hastings acceptance step.
@@
    For $l\in\{1,\ldots,L\}$, $(x_{l},m_{l})=\text{leapfrog}(x_{l-1},m_{l-1})$, as described in Algorithm F.1.
@@
Algorithm F.2 Hamiltonian Monte Carlo with stepsize $\epsilon>0$, mass matrix $M\in\mathbb{R}^{d\times d}$, and number of leapfrog steps $L\in\mathbb{N}$, starting from position $x\in\mathbb{R}^{d}$.
@@
F.2 Relation with MALA
@@
$\frac{\text{Normal}(x;x^{\prime}+(\epsilon^{2})/2M^{-1}\nabla\log\pi(x^{\prime}),\epsilon^{2}M^{-1})}{\text{Normal}(x^{\prime};x+(\epsilon^{2})/2M^{-1}\nabla\log\pi(x),\epsilon^{2}M^{-1})}.$ (F.1)
@@
The MALA acceptance ratio can be computed more efficiently than by separately computing two Normal density evaluations in (F.1), as described in Proposition 1 of Titsias (2024). Indeed the log-ratio can directly be computed as
$h(x,x^{\prime})-h(x^{\prime},x),\;\text{with}\;h(z,v)=\frac{1}{2}\left(z-v-\frac{\epsilon^{2}}{4}M^{-1}\nabla\log\pi(z)\right)^{T}\nabla\log\pi(v).$ (F.2)
F.3 Coupling of HMC
It is difficult to obtain exact meetings between two chains that evolve according to Hamiltonian Monte Carlo in general. Indeed, if the first chain at state $x$ draws a momentum $v$ and ends up in $x^{\prime}$, it is not obvious that a second chain at state $\tilde{x}$ could also arrive at state $x^{\prime}$ at the next step. Section of 4 Chen & Gatmiry (2023) addresses the question of existence of a momentum vector $\tilde{v}$ with which a leapfrog trajectory started at $\tilde{x}$ would end up in $x^{\prime}$. Even when such a momentum exists, it may not be available numerically. On the other hand, a simple coupling that consists of using the same momentum variable to propagate the two chains can result in a contraction (Mangoubi & Smith 2017, Bou-Rabee et al. 2020), i.e. the distance between the two chains goes to zero as the iterations go. Algorithm F.3 describes the common random numbers coupling of HMC.
@@ F.3 Coupling of HMC
    For $l\in\{1,\ldots,L\}$, $(x_{l},m_{l})=\text{leapfrog}(x_{l-1},m_{l-1})$, and $(\tilde{x}_{l},\tilde{m}_{l})=\text{leapfrog}(\tilde{x}_{l-1},\tilde{m}_{l-1})$, as described in Algorithm F.1.
@@
Algorithm F.3 Coupled Hamiltonian Monte Carlo with common random numbers, started from states $x$ and $\tilde{x}$.
In view of this Heng & Jacob (2019) propose the following strategy. Denote by $P_{\text{H}}$ the transition of HMC, and by $\bar{P}_{\text{H}}$ the common random number coupling of HMC. When chains are close to one another, a coupling $\bar{P}_{\text{M}}$ of an MRTH kernel $P_{\text{M}}$ with maximally coupled random walk proposals may result in an exact meeting (see Appendix A). Thus Heng & Jacob (2019) employ an MCMC algorithm with transition $\omega_{\text{H}}P_{\text{H}}+\omega_{\text{M}}P_{\text{M}}$, with $(\omega_{\text{H}},\omega_{\text{M}})$ probabilities summing to one, and they define its coupling as $\omega_{\text{H}}\bar{P}_{\text{H}}+\omega_{\text{M}}\bar{P}_{\text{M}}$. To generate these chains, with probability $\omega_{\text{H}}$ one employs the HMC kernel, otherwise one employs the MRTH kernel. The intuition is that the coupled HMC kernel brings the pair of chains closer to one another, and that the coupled MRTH kernel triggers exact meetings when the chains are close enough. The strategy is also used in Xu et al. (2021) for multinomial HMC.
In Section 5.2 we employ the following approach, suggested in the supplementary materials of Heng & Jacob (2019). We make the number of leapfrog steps $L$ random, and sample it uniformly in $\{1,\ldots,L_{\max}\}$. In the event $\{L=1\}$, HMC reverts to MALA, and we couple it with a reflection-maximal coupling of the proposal distributions, see Algorithms A.3 and F.4. When $\{L>1\}$ we employ a common random numbers coupling of HMC (Algorithm F.3). The tuning parameters are $\text{L}_{\max}$, $\epsilon$, and $M$. To summarise, the coupling of HMC employed in our experiments of Section 5.2 is in Algorithm F.5.
@@
Algorithm F.4 Coupled MALA with maximal coupling of proposals, started from $x$ and $\tilde{x}$.
@@
    If $L=1$, obtain $(x^{\prime},\tilde{x}^{\prime})$ from the coupled MALA transition in Algorithm F.4.
@@
    If $L>1$, obtain $(x^{\prime},\tilde{x}^{\prime})$ from the coupled HMC transition in Algorithm F.3, with $L$ leapfrog steps.
@@
Algorithm F.5 Coupling strategy for HMC, started from $x$ and $\tilde{x}$.
Appendix G MCMC for Bayesian linear regression with shrinkage prior
We provide details for the numerical experiments of Section 5.3. The example is taken from Biswas et al. (2022), where the authors consider couplings of the Gibbs sampler of Johndrow et al. (2020). We provide here a short and self-contained description of one particular version of the Gibbs sampler and its coupling; many more algorithmic considerations can be found in Biswas et al. (2022).
G.1 Model
@@ G.2.1 $\eta$-update
which we can target with the slice sampler described in Algorithm G.1, applied independently component-wise.
@@
Algorithm G.1 Iteration of slice sampling targeting $\eta_{j}\mapsto(\eta_{j}^{\frac{1-\nu}{2}}(1+\nu\eta_{j})^{\frac{\nu+1}{2}})^{-1}e^{-m_{j}\eta_{j}}$ on $(0,\infty)$.
G.2.2 $\xi$-update
@@ G.2.3 $\sigma^{2}$-update
G.2.4 $\beta$-update
With the notation $\Sigma=X^{T}X+\xi\text{Diag}(\eta)$, the distribution of $\beta$ given the rest is Normal with mean $\Sigma^{-1}X^{T}y$ and covariance matrix $\sigma^{2}\Sigma^{-1}$. We can sample from such Normals in a cost of order $n^{2}p$ using Algorithm G.2, as described in Bhattacharya et al. (2016).
@@ G.2.4 $\beta$-update
Algorithm G.1 Bhattacharya’s algorithm with common random numbers.
Algorithm G.2 Sampling from Normal$((X^{T}X+\xi\text{Diag}(\eta))^{-1}X^{T}y,\sigma^{2}(X^{T}X+\xi\text{Diag}(\eta))^{-1})$.
G.3 Coupled Gibbs sampler
We consider only one of the variants in Biswas et al. (2022), which is not necessarily the most efficient but achieves good performance in the experiments of Section 5.3 and is simpler than the “two-scale” coupling described in Biswas et al. (2022). We describe how to couple each update, with the first chain in state $\eta,\xi,\sigma^{2},\beta$ and the second in state $\tilde{\eta},\tilde{\xi},\tilde{\sigma}^{2},\tilde{\beta}$.
G.3.1 $\eta$-update
We consider two strategies to couple the slice sampling updates of $\eta_{j}$, as described in Algorithm G.1.
@@ G.3.1 $\eta$-update
    We can use a common uniform $V$ in the first step of Algorithm G.1, to define $U_{j}$ for the first chain and $\tilde{U}_{j}$ for the second. Then we can sample from a maximal coupling of the distributions of $\eta_{j}|U_{j}$ and $\tilde{\eta}_{j}|\tilde{U}_{j}$, using Algorithm A.2. This strategy results in a non-zero probability for the event $\{\eta_{j}=\tilde{\eta}_{j}\}$.
@@
    We can use a common uniform $V$ in the first step of Algorithm G.1, and then a common uniform $U^{*}$ in the third step. This is a pure “common random numbers” (CRN) strategy.

### v4

@@ Appendix D Verifying the assumptions in the AR(1) case
Algorithm D.1 Simulating pairs of ${\rm AR}(1,\phi,\sigma^{2})$ chains with reflection-maximal coupling.
@@
Lemma D.38.
For Algorithm D.1, the meeting time $\tau=\inf\{n\geq 1:X_{n}=Y_{n}\}$ satisfies
@@
We consider the following equivalent construction of $X$ and $Y$ as in Algorithm D.1. Let $(W_{n})$ be a sequence of independent $\text{Normal}(0,1)$ random variables. Set $X_{0}=x_{0}$ and define
@@
Combining Lemma D.38 with Proposition D.39, we obtain the following.
Corollary D.40.
For Algorithm D.1, the meeting time $\tau=\inf\{n\geq 1:X_{n}=Y_{n}\}$ satisfies
@@
where with the same constants as in Proposition D.39, $\tilde{C}(x,y)=\frac{2}{\tilde{\beta}}+|\frac{x-y}{2\sigma}|+3$ and $\bar{\beta}=\tilde{\beta}^{\frac{\log(\phi)}{\log(\tilde{\beta})+\log(\phi)}}$, and satisfies $\pi\otimes\pi(\tilde{C})<\infty$.
@@
We conclude by Lemma D.41. ∎
In combination with Lemma D.38, we obtain the following.
Corollary D.43.
For Algorithm D.1, the meeting time $\tau=\inf\{n\geq 1:X_{n}=Y_{n}\}$ satisfies
@@
where $\tilde{\beta}\in(0,1)$ is as in Proposition D.42, and $\tilde{C}(x,y)=1+(1-\phi^{2})\left(\frac{x_{0}-y_{0}}{2\sigma}\right)^{2}$ satisfies $\pi\otimes\pi(\tilde{C})<\infty$.
@@
where $\gamma=\beta^{\frac{\log(\alpha)}{\log(\beta)+\log(\alpha)}}$ and $\mathbb{P}_{\nu}(\tau>n)$ is as in Lemma D.44.
@@
Lemma D.44 provides that
@@
where we have used Devroye et al. (2018, Theorem 1.3) in the first line. From Proposition D.42, we find that
@@
we may deduce that $\mathbb{P}_{\mu}(\tau>n)\leq 2\tilde{\beta}^{n}$. Hence, we obtain by Corollary D.45,
@@
Appendix E Hamiltonian Monte Carlo
E.1 Basic version of HMC
Consider a target distribution on $\mathbb{R}^{d}$ with density $\pi$ and gradient of the log-density $\nabla\log\pi$, that can be evaluated pointwise. With Hamiltonian Monte Carlo (e.g. Neal 2011), we augment the state space $\mathbb{R}^{d}$ with a momentum variable, also taking values in $\mathbb{R}^{d}$. The momentum is periodically resampled from a $\text{Normal}(0,M)$ distribution, with covariance matrix $M\in\mathbb{R}^{d\times d}$. Algorithm E.1 describes one step of the leapfrog integrator, which updates a pair $(x,m)$ of position and momentum variables in a way that mimics certain Hamiltonian dynamics.
@@ E.1 Basic version of HMC
Algorithm E.1 Function $(x,m)\mapsto\text{leapfrog}(x,m)$, with stepsize $\epsilon>0$.
The transition kernel of HMC is described in Algorithm E.2. From an initial position, it draws an independent momentum from a centered multivariate Normal, then performs $L$ leapfrog steps, and finally the terminal position is accepted or not as the new state, with a Metropolis–Rosenbluth–Teller–Hastings acceptance step.
@@
    For $l\in\{1,\ldots,L\}$, $(x_{l},m_{l})=\text{leapfrog}(x_{l-1},m_{l-1})$, as described in Algorithm E.1.
@@
Algorithm E.2 Hamiltonian Monte Carlo with stepsize $\epsilon>0$, mass matrix $M\in\mathbb{R}^{d\times d}$, and number of leapfrog steps $L\in\mathbb{N}$, starting from position $x\in\mathbb{R}^{d}$.
@@
E.2 Relation with MALA
@@
$\frac{\text{Normal}(x;x^{\prime}+(\epsilon^{2})/2M^{-1}\nabla\log\pi(x^{\prime}),\epsilon^{2}M^{-1})}{\text{Normal}(x^{\prime};x+(\epsilon^{2})/2M^{-1}\nabla\log\pi(x),\epsilon^{2}M^{-1})}.$ (E.1)
@@
The MALA acceptance ratio can be computed more efficiently than by separately computing two Normal density evaluations in (E.1), as described in Proposition 1 of Titsias (2024). Indeed the log-ratio can directly be computed as
$h(x,x^{\prime})-h(x^{\prime},x),\;\text{with}\;h(z,v)=\frac{1}{2}\left(z-v-\frac{\epsilon^{2}}{4}M^{-1}\nabla\log\pi(z)\right)^{T}\nabla\log\pi(v).$ (E.2)
E.3 Coupling of HMC
It is difficult to obtain exact meetings between two chains that evolve according to Hamiltonian Monte Carlo in general. Indeed, if the first chain at state $x$ draws a momentum $v$ and ends up in $x^{\prime}$, it is not obvious that a second chain at state $\tilde{x}$ could also arrive at state $x^{\prime}$ at the next step. Section 4 of Chen & Gatmiry (2023) addresses the question of existence of a momentum vector $\tilde{v}$ with which a leapfrog trajectory started at $\tilde{x}$ would end up in $x^{\prime}$. Even when such a momentum exists, it may not be available numerically. On the other hand, a simple coupling that consists of using the same momentum variable to propagate the two chains can result in a contraction (Mangoubi & Smith 2017, Bou-Rabee et al. 2020), i.e. the distance between the two chains goes to zero as the iterations go. Algorithm E.3 describes the common random numbers coupling of HMC.
@@ E.3 Coupling of HMC
    For $l\in\{1,\ldots,L\}$, $(x_{l},m_{l})=\text{leapfrog}(x_{l-1},m_{l-1})$, and $(\tilde{x}_{l},\tilde{m}_{l})=\text{leapfrog}(\tilde{x}_{l-1},\tilde{m}_{l-1})$, as described in Algorithm E.1.
@@
Algorithm E.3 Coupled Hamiltonian Monte Carlo with common random numbers, started from states $x$ and $\tilde{x}$.
In view of this Heng & Jacob (2019) propose the following strategy. Denote by $P_{\text{H}}$ the transition of HMC, and by $\bar{P}_{\text{H}}$ the common random number coupling of HMC. When chains are close to one another, a coupling $\bar{P}_{\text{M}}$ of an MRTH kernel $P_{\text{M}}$ with maximally coupled random walk proposals may result in an exact meeting (see Appendix A.1). Thus Heng & Jacob (2019) employ an MCMC algorithm with transition $\omega_{\text{H}}P_{\text{H}}+\omega_{\text{M}}P_{\text{M}}$, with $(\omega_{\text{H}},\omega_{\text{M}})$ probabilities summing to one, and they define its coupling as $\omega_{\text{H}}\bar{P}_{\text{H}}+\omega_{\text{M}}\bar{P}_{\text{M}}$. To generate these chains, with probability $\omega_{\text{H}}$ one employs the HMC kernel, otherwise one employs the MRTH kernel. The intuition is that the coupled HMC kernel brings the pair of chains closer to one another, and that the coupled MRTH kernel triggers exact meetings when the chains are close enough. The strategy is also used in Xu et al. (2021) for multinomial HMC.
In Section 5.2 we employ the following approach, suggested in the supplementary materials of Heng & Jacob (2019). We make the number of leapfrog steps $L$ random, and sample it uniformly in $\{1,\ldots,L_{\max}\}$. In the event $\{L=1\}$, HMC reverts to MALA, and we couple it with a reflection-maximal coupling of the proposal distributions, see Algorithms A.3 and E.4. When $\{L>1\}$ we employ a common random numbers coupling of HMC (Algorithm E.3). The tuning parameters are $\text{L}_{\max}$, $\epsilon$, and $M$. To summarise, the coupling of HMC employed in our experiments of Section 5.2 is in Algorithm E.5.
@@
Algorithm E.4 Coupled MALA with maximal coupling of proposals, started from $x$ and $\tilde{x}$.
@@
    If $L=1$, obtain $(x^{\prime},\tilde{x}^{\prime})$ from the coupled MALA transition in Algorithm E.4.
@@
    If $L>1$, obtain $(x^{\prime},\tilde{x}^{\prime})$ from the coupled HMC transition in Algorithm E.3, with $L$ leapfrog steps.
@@
Algorithm E.5 Coupling strategy for HMC, started from $x$ and $\tilde{x}$.
Appendix F MCMC for Bayesian linear regression with shrinkage prior
We examine a Bayesian linear regression of $n=71$ responses on $p=4088$ predictors of the riboflavin data set (Bühlmann et al. 2014), using a continuous shrinkage prior on the coefficients (e.g. Bhadra et al. 2019). The model, MCMC algorithm and its coupling are taken from Johndrow et al. (2020), Biswas et al. (2022). Computational difficulties arise from the use of the shrinkage prior, that induces multimodality and heavy tails in the posterior distribution of the coefficients, denoted by $\beta$. The target is defined on the space of coefficients, global precision, local precisions, and the variance of the observation noise. The state space is of dimension $81782p+2=$ 8178. In this example, the meeting times have been shown to have Exponential tails in Biswas et al. (2022, Proposition 6) under the assumption that the global precision $\xi$ has a compact support, in which case Assumption 2 holds for any $\kappa$. In the experiments, we do not restrict the support of $\xi$, and we use a half-Cauchy distribution on $\xi^{-1/2}$.
We provide here a short and self-contained description of one particular version of the Gibbs sampler and its coupling; many more algorithmic considerations can be found in Biswas et al. (2022).
F.1 Model
@@ F.2.1 $\eta$-update
which we can target with the slice sampler described in Algorithm F.1, applied independently component-wise.
@@
Algorithm F.1 Iteration of slice sampling targeting $\eta_{j}\mapsto(\eta_{j}^{\frac{1-\nu}{2}}(1+\nu\eta_{j})^{\frac{\nu+1}{2}})^{-1}e^{-m_{j}\eta_{j}}$ on $(0,\infty)$.
F.2.2 $\xi$-update
@@ F.2.3 $\sigma^{2}$-update
F.2.4 $\beta$-update
With the notation $\Sigma=X^{T}X+\xi\text{Diag}(\eta)$, the distribution of $\beta$ given the rest is Normal with mean $\Sigma^{-1}X^{T}y$ and covariance matrix $\sigma^{2}\Sigma^{-1}$. We can sample from such Normals in a cost of order $n^{2}p$ using Algorithm F.2, as described in Bhattacharya et al. (2016).
@@ F.2.4 $\beta$-update
Algorithm F.1 Bhattacharya’s algorithm with common random numbers.
Algorithm F.2 Sampling from Normal$((X^{T}X+\xi\text{Diag}(\eta))^{-1}X^{T}y,\sigma^{2}(X^{T}X+\xi\text{Diag}(\eta))^{-1})$.
F.3 Coupled Gibbs sampler
We consider only one of the variants in Biswas et al. (2022), which is not necessarily the most efficient but achieves good performance in the experiments of Appendix F.4 and is simpler than the “two-scale” coupling described in Biswas et al. (2022). We describe how to couple each update, with the first chain in state $\eta,\xi,\sigma^{2},\beta$ and the second in state $\tilde{\eta},\tilde{\xi},\tilde{\sigma}^{2},\tilde{\beta}$.
F.3.1 $\eta$-update
We consider two strategies to couple the slice sampling updates of $\eta_{j}$, as described in Algorithm F.1.
@@ F.3.1 $\eta$-update
    We can use a common uniform $V$ in the first step of Algorithm F.1, to define $U_{j}$ for the first chain and $\tilde{U}_{j}$ for the second. Then we can sample from a maximal coupling of the distributions of $\eta_{j}|U_{j}$ and $\tilde{\eta}_{j}|\tilde{U}_{j}$, using Algorithm A.2. This strategy results in a non-zero probability for the event $\{\eta_{j}=\tilde{\eta}_{j}\}$.
@@
    We can use a common uniform $V$ in the first step of Algorithm F.1, and then a common uniform $U^{*}$ in the third step. This is a pure “common random numbers” (CRN) strategy.